OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x^2 - 7x - 9. Find f(g(-1)). Question 8 options: 1) -21 2) -3 3) 3 4) 21

OpenStudy (anonymous):

hello joemath

OpenStudy (anonymous):

yo :)

OpenStudy (anonymous):

guess what i had a huge "discussion" about two days ago

OpenStudy (anonymous):

what? o.O

OpenStudy (anonymous):

@marc \[g(-1)=-1\] \[f(-1)=-3\]

OpenStudy (anonymous):

here is a hint. what is \[\sqrt{16}\]?

OpenStudy (anonymous):

LOLOL

OpenStudy (anonymous):

First, work out what g(-1) is. In the equation \(g(x) = x^2-7x-9\), replace the x with -1.

OpenStudy (anonymous):

really it went on for ages

OpenStudy (anonymous):

you arent serious lolol

OpenStudy (anonymous):

guy tried to tell me that \[f(x)=\sqrt{x}\] is not a function

OpenStudy (anonymous):

LOLOL was it Chenna by any chance?

OpenStudy (anonymous):

unless you write \[f(x)=\sqrt{x}:x \geq0\]

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

i forget name. i almost gave up

OpenStudy (anonymous):

It was me!

OpenStudy (anonymous):

i guess people really dont understand square roots, that much >.< @Dalvoron was it really? o.O you dont seem like someone that would make that mistake.

OpenStudy (anonymous):

It's what textbooks tell me, it's what definitions tell me, it's what my lecturers have told me. I still believe it to be true, as I've found no evidence to the contrary, apart from \(\sqrt{x}\) referring only to the principal square root by convention in some cases, but not in others.

OpenStudy (anonymous):

There was a problem on here posted a while ago where we abused the fact that \[\sqrt{x} = |x|\] let me see if i can find it...

OpenStudy (anonymous):

then i have a question. why is quadatic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]? what is the plus or minus there for if the square root can be negative as well?

OpenStudy (anonymous):

mabye satellite can find it first, im looking for the piecewise function problem.

OpenStudy (anonymous):

i have yet to see a textbook that says \[\sqrt{a^2}=\pm a\]

OpenStudy (anonymous):

because if the square root also returned a negative value i dont think we could have solved it the way we did.

OpenStudy (anonymous):

or a text that said \[\sqrt{16}=\pm4\]

OpenStudy (anonymous):

Because the convention sometimes applies there I guess.

OpenStudy (anonymous):

have you seen this? http://www.wolframalpha.com/input/?i=sqrt%28x%29+%3D+-4

OpenStudy (anonymous):

convention "sometimes applies"? no the convention always applies. root means positive root. it is the inverse of the function \[f(x)=x^2: x \geq 0\]

OpenStudy (anonymous):

I have a text here (Text & Tests 1) which states "The square root of a number is the number which when multiplied by itself gives the number.", which would apply equally correctly to a positive or negative number.

OpenStudy (anonymous):

wow really? then i understand your confusion because if that is what it says, that is what it says. but if you look at the graph of \[y=\sqrt{x}\] you will see that it exists on the right hand side of the y axis, and only exists on theh left if you allow for complex numbers

OpenStudy (anonymous):

its not like i dont see where youre coming from, i was taught that the square root of a number is a plus or minus thing too. When i got to a analysis class though and got a problem wrong because of it though, i learned why thats not the case (in regards to functions, which is what that problem is about)

OpenStudy (anonymous):

For the sake of clarity, I can see why there should be universal acceptance of one or the other, but it just isn't the case. You probably see it more than I do as being only the principal square root, but that doesn't seem to be the case in Ireland.

OpenStudy (anonymous):

in fact i am completely baffled because when you derive the quadratic formula for example you start by saying that if \[x^2=9\] then \[x=\pm3\] and if \[x^2=5\] then \[x=\pm\sqrt{5}\]

OpenStudy (anonymous):

if the square root of 5 could be negative, then the plus minus bit is superfluous

OpenStudy (anonymous):

As much as I with maths always made sense to me, it doesn't :P Perhaps it's more widely accepted in the world of mathematics, but I studied physics mainly with some mathematics, and never saw that convention.

OpenStudy (anonymous):

i will believe you but i find it a bit bizarre that in Ireland the square root function is not well defined

OpenStudy (anonymous):

its because mathematics is taught sort of backwards imo. We use things before we truely understand them.

OpenStudy (anonymous):

i mean is certainly believe you since you say it. no reason to doubt you. could you send me a reference?

OpenStudy (anonymous):

Well when it's referred to as a function, naturally it can only refer to one of the roots, but we don't call it a function outside of the case where we specifically see \(f(x) = \sqrt{x}\), or some derivative of that.

OpenStudy (anonymous):

I mean, think about calculus. You start taking derivatives and integrals before you even learn the analytic side which lets you know when you are able to take derivatives and integrals.

OpenStudy (anonymous):

backwards!

OpenStudy (anonymous):

There was that textbook reference a few minutes ago. I'll try looking through some exam paper marking schemes where we would be penalised for not writing both roots.

OpenStudy (anonymous):

i checked here http://en.wikipedia.org/wiki/Square_root

OpenStudy (anonymous):

yes math is always taught backwards. annoying

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x). Question 7 options: 1) ± √x+81 2) ± 9√x 3) 1 over x^2-81 4) x^2 over 81

OpenStudy (anonymous):

and thats what creates this misunderstanding. You are taught one thing in middle/high school because its easier to grasp, and then taught something else in college.

OpenStudy (anonymous):

In the mean time: http://dictionary.reference.com/browse/square+root http://www.thefreedictionary.com/square+root Bear in mind, they're dictionaries, not mathematical textbooks of course :P

OpenStudy (anonymous):

It's unfortunate that I'm going to be a teacher, and I'm going to have to stick with the curriculum, rather than teaching what is "correct". I'd like to teach the use of \(\tau\) rather than \(\pi\) as well, but I can't do that either :(

OpenStudy (anonymous):

ahh i see. it seems to say that "square root" could be positive or negative, but that the symbol \[\sqrt{x}\] means the positive number

OpenStudy (anonymous):

That just makes it even more confusing.

OpenStudy (anonymous):

so i will admit to being incorrect about the term "square root" but correct in my assertion that \[\sqrt{x}\] is positive.

OpenStudy (anonymous):

thus for example 16 has two square roots, \[\sqrt{16}=4\] and \[-\sqrt{16}=-4\]

OpenStudy (anonymous):

and will have to stop myself from saying "plus or minus the square root of 16"

OpenStudy (anonymous):

im done >.< this is the second time this has come up ( with the same problem no less! lol)

OpenStudy (anonymous):

With any luck, that will propagate over here, and the whole issue will not exist in future.

OpenStudy (anonymous):

i wonder what i can say instead?

OpenStudy (anonymous):

Refer only to the principal square root :P

OpenStudy (anonymous):

btw not to beat a dead horse, but the question where this came up had to do with whether \[\sqrt{10}\] was greater than 3. i maintain that it is greater than three because the root symbol means positive one

OpenStudy (anonymous):

in other words i maintain that \[\sqrt{10}\] is not some number whose square is ten, but the positive number whose square is ten

OpenStudy (anonymous):

Fair enough. For the purposes of this site at least, I'll try to make the distinction

OpenStudy (anonymous):

beat that dead horse! lol

OpenStudy (anonymous):

glad that is cleared up, at least for me

OpenStudy (anonymous):

someone give me a medal so i can be lvl 60 lol >.>

OpenStudy (anonymous):

nvm, thanks dalvoron :) and satellite :P

OpenStudy (anonymous):

Congrats on hitting 60!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

@dalvoron am sure you will make an excellent math teacher, and enjoy it too. outside of the vow of poverty you will have to take.

OpenStudy (anonymous):

just please don't ask your students ever to "simplify" please please please.

OpenStudy (anonymous):

i couldnt fine the piecewise problem :( forever lost!

OpenStudy (anonymous):

I'm caught between a rock and a hard place. Ideally, I should teach them about the truth of mathematics, and be as clear as possible; but at the end of the day, they have to be prepared for the exams that come up. Perhaps an educational reform is in order, and I can squeeze in my love of \(\tau\) as well!

OpenStudy (anonymous):

do it in secret <.< a secret society of truth!

OpenStudy (anonymous):

I could totally do that. Conspire with my fellow maths teachers as well. Write a textbook with subtle references to the truth.

OpenStudy (anonymous):

i think i saw it here, maybe, an otherwise excellent video, where he says "simplify" to mean three completely different things. sometimes it means add to get 0, sometimes it means cancel to get one, and sometimes it means that a function composed with its inverse "simplfies" so give what you started with

OpenStudy (anonymous):

Every now and then, it becomes apparent that we humans can be really stupid.

OpenStudy (anonymous):

if you want a laugh look at the implicit differentiation video here http://justmathtutoring.com/

OpenStudy (anonymous):

I was never great at implicit differentiation, but I think I just got worse.

OpenStudy (anonymous):

Let f(x) = 8x^3 - 22x^2 - 4 and g(x) = 4x - 3. Question 4 options: 1) 2x^2 - 4x - 3 - (13 over 4x-3) 2) 2x^2 - 4x - 3 - (4x-3 over 13) 3) 2x^2 - 7x - 1 4) 2x^2 - 7x - 5 + (x-4 over 4x-3)

OpenStudy (anonymous):

marc you should repost this question anew. no one will find it hear. it is also incomplete because we are not sure that to do with f and g

OpenStudy (anonymous):

Sorry about highjacking the question, but I think there are some helpful answers in there somewhere.